OpenStudy (anonymous):

Find the equation of the line passing through point (4,2) and perpendicular to y=1/3 X - 7

OpenStudy (anonymous):

y-b=m(x-a) y-2=-3(x-4) y=-3x+14 The gradient is -3 as 1/3 * -3 = -1